# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

### v2

@@ Abstract
Satellites are increasingly adopting on-board AI for enhanced autonomy through in-orbit inference. In this context, the use of deep learning (DL) techniques for segmentation in hyperspectral (HS) satellite imagery offers advantages for remote sensing applications, and therefore, we train 16 different models, whose codes are made available through our study, which we consider to be relevant for on-board multi-class segmentation of HS imagery, focusing on classifying oceanic (sea), terrestrial (land), and cloud formations. We employ the HYPSO-1 mission as an illustrative case for sea-land-cloud segmentation, and to demonstrate the utility of the segments, we introduce a novel sea-land-cloud ranking application scenario. We consider how to prioritize HS image downlink based on sea, land, and cloud coverage levels from the segmented images. We comparatively evaluate the models for future in-orbit deployment, considering performance, parameter count, and inference time. The models include both shallow and deep models, and after we propose four new DL models, we demonstrate that segmenting single spectral signatures (1D) outperforms 3D data processing comprising both spectral (1D) and spatial (2D) contexts. We conclude that our lightweight DL model, called 1D-Justo-LiuNet, consistently surpasses state-of-the-art models for sea-land-cloud segmentation, such as U-Net and its variations, in terms of performance (0.93 accuracy) and parameter count (4,563). However, the 1D models present longer inference time (15s) in the tested processing architecture, which seems to be a suboptimal architecture for this purpose. Finally, after demonstrating that in-orbit segmentation should occur post L1b radiance calibration rather than on raw data, we also show that reducing spectral channels down to 3 lowers models’ parameter counts and inference time, at the cost of weaker segmentation performance.
keywords:
@@
Remote Sensing , Satellite Imagery , Hyperspectral Imaging , Deep Learning , Segmentation
[…]
[inst1]organization=Dept. Engineering Cybernetics, Norwegian University of Science and Technology,city=Trondheim, country=Norway,
[…]
[inst2]organization=Dept. of Computer Science, University of Bucharest,country=Romania,
[…]
[inst3]organization=Independent Consultant,city=Montreal, country=Canada,
1 Introduction
@@
Artificial Intelligence (AI) has seen a growing utilization aboard satellite platforms for remote sensing, benefiting Earth observation missions equipped with various technologies from conventional RGB to both Multispectral Imaging (MSI) and Hyperspectral Imaging (HSI). The importance of monitoring the Earth’s environment with HSI technology is emphasized by the use of hyperspectral (HS) imagers in numerous contemporary satellite missions, e.g., ESA/Phi-Sat-1 [1], ASI/PRISMA [2, 3], DLR/EnMAP [4], and NTNU/HYPSO-1 [5]. The importance of HS imagers on satellites is further highlighted by upcoming missions, e.g., ESA/CHIME [6, 7], NASA/PACE, GLIMR and SBG [8], ASI/PRISMA Second Gen [9], as well as NTNU/HYPSO-2 and NTNU/HYPSO-3 [10]. HS missions offer great potential due to their detailed spectral characterization of Earth’s features, but they also share a common limitation with other optical remote sensing technologies: the challenge of cloud formation in the atmosphere obstructing the satellites’ capacity to monitor targets of interest. In this context, techniques such as semantic segmentation proves valuable in a broad range of space applications, e.g., from cloud recognition to wildfire detection [11, 12, 13, 14, 15, 16, 17, 18]. Segmentation can indeed serve as a cloud detection technique, categorizing image pixels as either cloudy or non-cloudy, essentially performing binary classification. As a mission example, after the launch in 2020 of the Phi-Sat-1 HS mission part of ESA’s EOP initiative [1], a significant milestone was marked in 2021 when, for the first time, on-board AI was introduced in a satellite for object detection, with the initial focus being the detection of clouds aboard the satellite. Phi-Sat-1 marked the first in-orbit inference with its convolutional CloudScout Deep Learning (DL) model for cloud detection [11, 14]. The mission’s imager, HyperScout-2, captures the Visible and Near-Infrared (VNIR) and Thermal Infrared (TIR) spectral ranges, and subsequently deploys data processing via AI models in its on-board configurable platform facilitating updates for its algorithms in flight [1, 12]. These in-orbit updates rely on an embedded configurable Commercial off-the-shelf (COTS) platform, which is based on a Vision Processing Unit (VPU) [12, 19]. Furthermore, beyond the Phi-Sat-1 mission, the literature [13] also extends cloud segmentation to image data of the missions Sentinel-2B (MSI) and FACSAT-1 (RGB) [20]. For instance, the state of the art [13] deploys a satellite data priority system to compute the cloud coverage level via DL-based segmentation on microcontrollers (MCUs), which determines the image downlink priority. Furthermore, the literature demonstrates that cloud detection can enhance data management aboard satellite platforms as presented next. On the one hand, selective HSI compression methods may be applied to cloud segments via CCSDS-123.0-B-2 to reduce their encoding rate [21, 22], while other methods like Compressed Sensing (CS) can also be employed to handle cloudy data, where CS proves particularly advantageous for satellites as it transfers computational complexity from orbit to mere ground-based reconstruction [23, 24]. Therefore, segmentation can lead to selective compression reducing the bit rate of data segregated as cloudy, or its absolute removal by data screening [12]. Consequently, this streamlines data postprocessing and enables the satellite to downlink, for instance, only valuable non-cloudy data, thus optimizing the usage of the often limited communication channel [13]. On the other hand, on-board cloud segmentation can additionally play a role in reducing decision-making latency, increasing satellite autonomy with more optimal on-board operations [12, 13, 15, 25, 26, 27, 28]. Furthermore, from a broader perspective, the models first trained for cloud binary segmentation can be trained again to solve more complex problems instead, e.g., multi-class segmentation, where each image pixel is assigned to one of various categories. To further illustrate the importance of AI-powered cloud and multi-class segmentation for space applications, we next employ a particular small satellite mission as an example case. Namely, for Earth observation initiatives such as HYPSO-1 (HYPerspectral small Satellite for Ocean observation) [5], the reduction in decision-making latency is especially important for time-sensitive remote sensing applications, such as early event warnings. HYPSO-1 focuses on timely detection of marine phenomena such as Harmful Algal Blooms (HABs), which can lead to significant mortality of farmed salmon. To this extent, in-orbit segmentation for adequate on-board data management can play a crucial role in achieving prompt management responses to mitigate adverse effects in such scenarios. In the context of multi-class segmentation, by segregating the data along various categories of interest, on-board AI can activate other subsequent agents in observational hierarchies monitoring the environment at multiple spatial, spectral and temporal scales with e.g. drones and marine vessels [5, 29, 30], where the HYPSO-1 satellite may serve as the space segment of an observational pyramid. However, handling increasingly advanced tasks on satellites, such as segmentation of growing complexity, is feasible only when the COTS components within the satellite’s payload are appropriately designed and optimized. Some COTS platforms provide benefits by accelerating, with high power efficiency, the repetitive operations commonly found in DL networks [1, 14, 19, 25]. Indeed, to enable AI applications in satellites, the literature explores various COTS hardware from the conventional MCUs [13] to the high-performance VPUs aboard Phi-Sat-1, Field-Programmable Gate Arrays (FPGAs) on board HYPSO-1, Tensor Processing Units (TPUs), Graphics Processing Units (GPUs), and other dedicated accelerators [1, 10, 11, 31, 32, 33].
@@
To our knowledge, the existing literature in the field of cloud detection on board satellite platforms does not compare how different shallow and deep learning models, including distinctions like 1D vs. 2D-CNN-based models, influence the performance and complexity of cloud segmentation. Furthermore, the field also overlooks the effects of reducing the number of spectral channels, particularly in MSI/HSI data, on models for cloud segmentation. The existing literature lacks comparisons of key aspects, such as whether on-board AI for cloud detection, is better applied to raw data or L1b calibrated radiance data. It does not discuss how L1b calibration impacts segmentation results and does not provide justification for the optimal stage in the on-board data processing pipeline for image segmentation. Moreover, within the literature pertaining to cloud detection in HS satellite imagery, the primary emphasis naturally lies on binary segmentation, specifically distinguishing between the presence of clouds and non-cloudy data; in fact, as earlier mentioned, there are ranking systems that assess cloud levels based on the segmented outcomes. Nevertheless, none of the systems at the time of writing perform multi-class semantic segmentation for the simultaneous detection of clouds, sea, and land categories, which typically account for most of the variance in HS satellite imagery. The absence of sea-land-cloud segmentation systems impedes, as a result, the development of applications like priority ranking systems that can integrate the three categories into a unified ranker. In this work, we tackle the mentioned research gaps and utilize the HYPSO-1 mission as a case study, using its “Sea-Land-Cloud-Labeled Dataset” openly accessible in our prior article [34] (see data availability section at the end of this paper).
@@
Following a methodology inspired by the data science CRISP-DM approach (CRoss Industry Standard Process for Data Mining) to ensure robust and reliable results, we train and compare 16 different Machine Learning (ML) and DL models. This extends the segmentation model baseline for HYPSO-1 first established in [34], and our research findings make contributions to the field as follows:
  • 1.
@@
    We propose and fine-tune four new 1D-CNN and 2D-CNN models, with the aim of adapting them for on-board deployment on satellites. This optimization entails reducing the number of model parameters to achieve highly compact lightweight models while maintaining performance.
  • 2.
@@
    For the case context of multi-class sea-land-cloud HS segmentation for the HYPSO-1 mission, we demonstrate that 1D-CNN models exhibit superior performance when compared to 2D-CNNs. Indeed, our optimized 1D-CNN lightweight models, 1D-Justo-LiuNet and 1D-Justo-HuNet, consistently outperform other state-of-the-art models.
  • 3.
@@
    We substantiate that segmenting L1b radiance data, as opposed to using raw sensor data, enhances the segmentation of the top-performing models.
  • 4.
@@
2 Methodological Background
2.1 Data Source: A Hyperspectral Labeled Datastet from HYPSO-1
@@
We employ the “Sea-Land-Cloud-Labeled Dataset” from HYPSO-1 [34] (see data availability section). The dataset consists of pixel-level labels for 25 million pixels across 38 scenes captured by HYPSO-1, where each HS image has spatial dimension of 956 $\times$ 684 pixels. Each pixel, contains a spectral signature of 120 channels covering the 400 to 800 nm range in the VNIR spectrum, with approximately 5 nm spectral resolution [5, 29]. The dataset consists of diverse scenes from various environments and provides raw data and L1b calibrated radiance for comparison with other HS missions. The dataset includes expert pixel-level annotations for three classes: “sea”, “land”, and “clouds”. The dataset exhibits a slight class imbalance with 37.01% pixels representing sea, 40.14% representing land, and 22.85% representing clouds/overexposed areas.
2.2 Channel Reduction Preprocessing
Following the common practice to inspect the data features prior to training ML and DL models, we next identify any unusual channel features that may require correction. In the sea-land-cloud dataset [34], we observe that the first four channels with wavelengths $\lambda$=388, 391, 395, and 399 nm (visible blue spectrum) exhibit data samples equal to 0 for all the dataset HS images. We observe that this phenomenon occurs only to the L1b radiance, however, it is not present in the raw sensor data. This suggests that the occurrence of zero values in the first channels is a result of the radiometric calibration procedures applied by [34] following [35]. Furthermore, we observe that the dataset, near its wavelength $\lambda$=760 nm (oxygen A-band within the NIR spectrum), presents a sharp decline in light intensity which is attributed to the significant light absorption by the atmosphere’s oxygen at this wavelength [36, 37]. In summary, we exclude first wavelengths $\lambda$=388, 391, 395, and 399 nm, and we also exclude the wavelengths in close proximity to $\lambda$=760 nm, namely, $\lambda$=758, 761, 765, and 768 nm. As a result, the dataset’s original 120 channels are reduced to 112. The 112 channels, after undergoing min-max normalization, are the ones we employ for training the ML and DL models in the subsequent methodology. Since this work additionally studies the impact on the segmentation outcomes when significantly reducing the number of channels for the particular context of sea-land-cloud HSI segmentation, we use the 112-channel data to create a new dataset containing only 3 channels as detailed in A.
2.3 Segmentation Modeling
2.3.1 Processing of 1D vs 3D Data
@@
Our method involves training 16 distinct ML and DL models that are considered of interest for in-orbit deployment, with a particular emphasis on DL models, as they are often better suited to capture more complex data relationships. To train the models, we maintain the same data split used in our prior article [34], allocating 30 HS images for training (79%), 3 images for validation (8%), and 5 images for testing (13%). Furthermore, we employ a deployment set of 30 additional HS images, also from the sea-land-cloud dataset, to test the best-performing model’s ability to generalize to further new unseen data. Additional details such as the specific image identifiers for the data split are provided in the supplementary material (see data availability section).

### v3

# Semantic Segmentation in Satellite Hyperspectral Imagery by Deep Learning

@@ Abstract
Satellites are increasingly adopting on-board AI to optimize operations and increase autonomy through in-orbit inference. The use of Deep Learning (DL) models for segmentation in hyperspectral imagery offers advantages for remote sensing applications. In this work, we train and test 20 models for multi-class segmentation in hyperspectral imagery, selected for their potential in future space deployment. These models include 1D and 2D Convolutional Neural Networks (CNNs) and the latest vision transformers (ViTs). We propose a lightweight 1D-CNN model, 1D-Justo-LiuNet, which outperforms state-of-the-art models in the hypespectral domain. 1D-Justo-LiuNet exceeds the performance of 2D-CNN UNets and outperforms Apple’s lightweight vision transformers designed for mobile inference. 1D-Justo-LiuNet achieves the highest accuracy (0.93) with the smallest model size (4,563 parameters) among all tested models, while maintaining fast inference. Unlike 2D-CNNs and ViTs, which encode both spectral and spatial information, 1D-Justo-LiuNet focuses solely on the rich spectral features in hyperspectral data, benefitting from the high-dimensional feature space. Our findings are validated across various satellite datasets, with the HYPSO-1 mission serving as the primary case study for sea, land, and cloud segmentation. We further confirm our conclusions through generalization tests on other hyperspectral missions, such as NASA’s EO-1. Based on its superior performance and compact size, we conclude that 1D-Justo-LiuNet is highly suitable for in-orbit deployment, providing an effective solution for optimizing and automating satellite operations at edge.
Index Terms:
@@
Remote Sensing, Satellite Hyperspectral Imagery, Segmentation, Deep Learning, 1D-CNNs, 2D-CNNs, ViTs.
I Introduction
@@
Artificial Intelligence (AI) has seen a growing utilization aboard satellite platforms for remote sensing, benefiting Earth observation missions equipped with various technologies, from conventional RGB to Multispectral Imaging (MSI) and Hyperspectral Imaging (HSI). The importance of monitoring the Earth’s environment with HSI technology is emphasized by the use of hyperspectral (HS) imagers in numerous contemporary satellite missions. Examples include $\Phi$-sat missions from the European Space Agency (ESA) [1], PRecursore IperSpettrale della Missione Applicativa (PRISMA) from the Italian Space Agency (ASI) [2, 3], the Environmental Mapping and Analysis Program (EnMAP) from the German Aerospace Center (DLR) [4], and the HYPerspectral small Satellite for Ocean observation (HYPSO) missions from the Norwegian University of Science and Technology (NTNU) [5], with HYPSO-1 launched in 2022 and HYPSO-2 in August 2024. The importance of HS imagers on satellites is further highlighted by upcoming missions such as the The Copernicus Hyperspectral Imaging Mission for the Environment (CHIME) from ESA [6, 7], the Plankton, Aerosol, Cloud, ocean Ecosystem (PACE), the Geosynchronous Littoral Imaging and Monitoring Radiometer (GLIMR), the Surface Biology and Geology (SBG) from the National Aeronautics and Space Administration (NASA) [8], ASI/PRISMA Second Gen [9], as well as NTNU/HYPSO-3 [10]. HS missions offer great potential due to their detailed spectral characterization of Earth’s features, but they also share a common limitation with other optical remote sensing technologies: the challenge of cloud formation in the atmosphere obstructing the satellites’ capacity to monitor targets of interest on the Earth’s surface. In this context, techniques such as semantic segmentation proves valuable in a broad range of space applications, e.g., from cloud recognition to wildfire detection [11, 12, 13, 14, 15, 16, 17, 18]. Segmentation can indeed serve as a cloud detection technique, categorizing image pixels as either cloudy or non-cloudy, essentially performing binary classification. As a mission example, after the launch in 2020 of the $\Phi$-sat-1 HS mission part of ESA’s Earth Observation Programmes [1], a significant milestone was marked in 2021 when, for the first time, on-board AI was introduced in a satellite for object detection, with the initial focus being the detection of clouds aboard the satellite. $\Phi$-sat-1 marked the first in-orbit inference at edge with its convolutional CloudScout Deep Learning (DL) model for cloud detection [11, 14]. The mission’s imager, HyperScout-2, captures the Visible and Near-Infrared (VNIR) and Thermal Infrared (TIR) spectral ranges, and subsequently deploys data processing via AI models in its on-board configurable platform facilitating updates for its algorithms in flight [1, 12]. These in-orbit updates rely on an embedded configurable Commercial Off-The-Shelf (COTS) platform, which is based on a Vision Processing Unit (VPU) [12, 19]. Furthermore, beyond the $\Phi$-sat-1 mission, the literature [13] also extends cloud segmentation to image data of the missions Sentinel-2B (MSI) and FACSAT-1 (RGB) [20]. For instance, the state of the art [13] deploys a satellite data priority system to compute the cloud coverage level via DL-based segmentation on microcontrollers (MCUs), which determines the image downlink priority. Furthermore, the literature demonstrates that cloud detection can enhance data management aboard satellite platforms as presented next. First, selective HSI compression methods may be applied to cloud segments via CCSDS-123.0-B-2 to reduce their encoding rate [21, 22], while other methods such as Compressed Sensing (CS) can also be employed to handle cloudy data, where CS proves particularly advantageous for satellites, as it transfers computational complexity from orbit to ground-based reconstruction [23, 24]. Therefore, segmentation can lead to selective compression reducing the bit rate of data segregated as cloudy, or its absolute removal by data screening [12]. This enables the satellite to downlink, for instance, only valuable non-cloudy data, thus optimizing the usage of the often limited communication channel [13]. On-board cloud segmentation can also play a role in reducing decision-making latency, increasing satellite autonomy with more optimal operations [12, 13, 15, 25, 26, 27, 28]. Furthermore, from a broader perspective, the models first trained for cloud binary segmentation can be retrained to solve more complex problems instead, e.g., multi-class segmentation, where each image pixel is assigned to one of various categories. To further illustrate the importance of AI-driven cloud and multi-class segmentation for space applications, we next employ a particular small satellite mission as a case study. In this work, we employ the Earth observation CubeSat HYPSO-1 [5]. In this mission, reducing decision-making latency is especially important for time-sensitive remote sensing applications, such as early event warnings. HYPSO-1 focuses on timely detection of marine phenomena such as Harmful Algal Blooms (HABs), which can lead to significant mortality of farmed salmon. In-orbit segmentation for adequate on-board data management can play a crucial role enabling prompt management responses to mitigate adverse effects in such scenarios. In the context of multi-class segmentation, by segregating the data along various categories of interest, on-board AI can activate other subsequent agents in observational hierarchies monitoring the environment at multiple spatial, spectral and temporal scales with e.g. drones and marine vessels [5, 29, 30], where the HYPSO-1 satellite may serve as the space segment of an observational pyramid. However, handling increasingly advanced tasks on satellites, such as segmentation of growing complexity, is feasible only when the COTS components within the satellite’s payload are appropriately designed and optimized. Some COTS platforms provide benefits by accelerating, with high power efficiency, the repetitive operations commonly found in DL networks [1, 14, 19, 25]. Indeed, to enable AI applications in satellites, the literature explores various COTS hardware, from the conventional MCUs [13] to the high-performance VPUs aboard $\Phi$-sat-1, Field-Programmable Gate Arrays (FPGAs) on board HYPSO missions, Tensor Processing Units (TPUs), Graphics Processing Units (GPUs), amog other dedicated accelerators [1, 10, 11, 31, 32, 33].
@@
Despite the extensive use of 2D Convolutional Neural Networks (CNNs) in state-of-the-art semantic segmentation for hyperspectral imaging, our primary research question is: How can we compress deep models to create lighter versions for future in-orbit deployment, while not only maintaining but also improving performance compared with state-of-the-art methods, from standard UNet architectures to the latest advanced vision transformers? Furthermore, to the best of our knowledge, the existing literature in the field of cloud detection on board satellite platforms does not compare how different shallow and deep learning models, including distinctions such as 1D vs. 2D-CNN-based models and Vision Transformers (ViT), influence the performance and complexity of cloud segmentation. Additionally, the field also overlooks the effects of reducing the number of spectral channels, particularly in MSI/HSI data, on models for cloud segmentation. The existing literature lacks comparisons of key aspects, such as whether on-board AI for cloud detection, is better applied to raw data or calibrated data. It does not discuss how L1b calibration impacts segmentation results and does not provide justification for the optimal stage in the on-board data pipeline for image segmentation at edge. Moreover, within the literature pertaining to cloud detection in HS satellite imagery, the primary emphasis naturally lies on binary segmentation, distinguishing between the presence of clouds and non-cloudy data; in fact, as earlier mentioned, there are ranking systems that assess cloud levels based on the segmented outcomes. Nevertheless, none of the systems at the time of writing perform multi-class semantic segmentation for the simultaneous detection of clouds, sea, and land categories, which typically account for most of the variance in HS satellite data. The absence of sea-land-cloud segmentation systems impedes, as a result, the development of applications such as priority ranking systems that can integrate the three categories into a unified ranker. In this work, we tackle the mentioned research gaps and utilize the HYPSO-1 mission as a case study, using its “Sea-Land-Cloud-Labeled Dataset” openly accessible in our prior article [34] (see data availability section at the end of this paper).
@@
Following a methodology inspired by the data science CRISP-DM approach (CRoss Industry Standard Process for Data Mining) to ensure robust and reliable results, we train and compare 19 different machine learning (ML) models. For further comparison, we will train an additional model, bringing the total to 20 models in this study. This extends the segmentation model baseline for HYPSO-1 first established in [34], and our research findings make contributions to the field as follows:
@@
  • We propose and fine-tune four new 1D-CNN and 2D-CNN models, with the aim of adapting them for on-board deployment on satellites. This optimization entails reducing the number of model parameters to achieve highly compact lightweight models, while maintaining performance.
@@
  • In the context of multi-class HS segmentation, we demonstrate that 1D-CNN models exhibit superior performance when compared with 2D-CNN UNets and FastViT. Indeed, our lightweight 1D-CNN models, 1D-Justo-LiuNet and 1D-Justo-HuNet, consistently outperform the state of the art. Our results are confirmed by generalization tests on different satellites.
@@
  • We substantiate that segmenting L1b radiance data, as opposed to using raw sensor data, slightly enhances the segmentation of top-performing models.
@@
II Methodological Background
II-A Datasets: Labeled Satellite Imagery
@@
We perform experiments on three datasets. The first one is the Sea-Land-Cloud-Labeled Dataset from HYPSO-1 [34] [^1] https://ntnu-smallsat-lab.github.io/hypso1_sea_land_clouds_dataset/. The dataset consists of pixel-level labels for 25 million pixels across 38 scenes captured by HYPSO-1, where each HS image has a spatial dimension of 956 $\times$ 684 pixels. Each pixel contains a spectral signature of 120 channels covering the 400 to 800 nm range in the visible and VNIR spectrum, with an approximate spectral resolution of 5 nm and a spatial resolution of approximately 100 m $\times$ 600 m in the across- and along-track directions [5, 29]. The dataset consists of diverse scenes from various environments and provides raw data and L1b calibrated radiance for comparison with other satellite missions. The dataset includes expert pixel-level annotations for three classes: sea, land, and clouds. The data exhibits a slight class imbalance with 37.01% pixels representing sea, 40.14% representing land, and 22.85% representing clouds/overexposed areas.
@@
Our method involves training 19 distinct models that are considered of interest for in-orbit deployment, with a particular emphasis on DL models, as they are often better suited to capture more complex data relationships. To train the models, we maintain the same data split used in our previous work [34], allocating 30 HS images for training (79%), 3 images for validation (8%), and 5 images for testing (13%). Furthermore, we employ a deployment set of 30 additional HS images, from the same sea-land-cloud dataset, to test the models ability to generalize to further new unseen data. In summary, we test on 35 different scenes in total from the deployment and test sets. The scenes comprise coastal zones, forests, dry regions, cloudy areas, and more. The deployment set is provided in the supplementary material of this article (see data availability section), where additional details, such as the specific image identifiers for the data split, are also included.
Furthermore, to assess the generalization of the models in other satellites, we test them on two additional datasets. First, we perform experiments using 35 calibrated HS images from NASA’s Earth Observing-1 (EO-1) mission [35] [^2] https://earthexplorer.usgs.gov/. Second, we conduct experiments on Satellite Imagery of Dubai (SID) dataset [^3] https://humansintheloop.org/resources/datasets/semantic-segmentation-dataset-2/. For EO-1, the mission utilizes the Hyperion sensor [36], delivering high spatial (30 m) and spectral (10 nm) resolutions. Our EO-1 dataset includes expert ground-truth annotations with labels: no cloud, thin cloud, and thick cloud. Further details on this dataset can be found in our article [37]. In the case of the SID dataset, it comprises 72 RGB aerial images of Dubai gathered by the Mohammed Bin Rashid Space Center (MBRSC). The images come with annotations at the pixel level with 6 class labels: buildings, roads, land, vegetation, water, and other.
II-B Channel Reduction Preprocessing
To begin, we emphasize that the main goal of our work is not to find the optimal dimensionality reduction for our application. Instead, we aim to identify any features in the data that could impact our models and, more specifically, to examine how segmentation results are influenced when dimensionality is significantly reduced.
For data preprocessing, we first identify in the HYPSO-1 dataset unusual channel features that may require reduction. In the sea-land-cloud dataset [34], we observe that the first four channels with wavelengths $\lambda$=388, 391, 395, and 399 nm (visible blue spectrum) exhibit data samples equal to 0 for all the dataset images. We observe that this phenomenon occurs only for the L1b radiance, however, it is not present in the raw sensor data. This suggests that the occurrence of zero values in the first channels is a result of the radiometric calibration procedures applied by [34], following [38]. Furthermore, we observe that the dataset, near its wavelength $\lambda$=760 nm (oxygen A-band within the NIR spectrum), presents a sharp decline in light intensity which is attributed to the significant light absorption by the atmosphere’s oxygen at this wavelength [39, 40]. In summary, we exclude first wavelengths $\lambda$=388, 391, 395, and 399 nm, and we also exclude the wavelengths in close proximity to $\lambda$=760 nm, namely, $\lambda$=758, 761, 765, and 768 nm. As a result, the dataset’s original 120 channels are reduced to 112. The 112 channels, after undergoing min-max normalization, are the ones we employ for training the ML and DL models in the subsequent methodology. Furthermore, this work also examines the impact of significantly reducing the number of channels on sea-land-cloud HSI segmentation. Therefore, we use the 112-channel data to create a new dataset containing only 3 channels, since the legacy algorithms for cloud detection in HS satellite imagery often select 3 channels to mimic RGB image processing [11, 12, 13, 14]. Details about the dimensionality reduction algorithm used, which is beyond the scope of this work, can be found in Appendix A. In the appendix, we explain that the 3 channels selected correspond to 412.72 nm for the blue spectrum, 699.61 nm for the green and red spectra, and 747.77 nm for the NIR spectrum.
II-C Semantic Segmentation Modeling
II-C1 Processing of 1D vs. 3D Data